OpenStudy (anonymous):

4_06]Choose the equation below that represents the line passing through the point (2, -5) with a slope of -3. y = -3x - 13 y = -3x + 11 y = -3x + 13 y = -3x + 1

OpenStudy (anonymous):

need help

OpenStudy (anonymous):

plz

OpenStudy (anonymous):

y = -3x + 1 -5 = -3*2 +1 -5 = -5